A large fast-food restaurant is having a promotional game where game pieces can be found on various products. Customers can win food or cash prizes. According to the company, the probability of winning a prize (large or small) with any eligible purchase is 0.15.
Consider your next 20 purchases that produce a game piece. Calculate the following:
What is the probability that you win between 4 and 6 (inclusive) prizes?
This is a binomial distribution. Give your answers to 4 decimal places.

Answer by Boreal(15235)

p(4)=20C4*0.15^4*0.85^16=0.1821
p(5)=20C5*0.15^5*0.85^15=0.1028
p(6)=20C6*0.15^6*0.85^14=0.0454
=sum or 0.3303